Mary and Janet are running in the Apple Hill Run. Mary runs at 7mph, Janet at 5 mph. If they start at the same time, how long will it be before they are 1.5 miles apart?
Answer by stanbon(75887)

Mary and Janet are running in the Apple Hill Run. Mary runs at 7mph, Janet at 5 mph. If they start at the same time, how long will it be before they are 1.5 miles apart?
--------------------
Mary DATA:
rate = 7 mph ; time = t hrs. ; distance = 7t miles
-----
Janet DATA:
rate = 5 mph ; time = t hrs. ; distance = 5t miles
-----
Equation:
7t-5t = 1.5 hrs
2t = 1.5 hrs
time = (3/4)hr = 45 minutes
=================
Cheers,
Stan H.
==============